OpenStudy (anonymous):

In oder for a satellite to remain in a circular orbit above the same spot on Earth ,the satellite must be 35,800 kilometers above the equator. Write an equation for the orbit of the satellite .Use the center of Earth as the origin and 6400 kilometers for the radius of Earth.

OpenStudy (anonymous):

Taking (0,0) as the center of the earth, and of the orbit. General cartesian eqn for the circle : (x-a)^2 + (y-b^)2 = r^2 But (a,b) is (0,0), as stated above So : x^2 + y^2 = r^2 What is r ? r = 6400 + 35800 = 42200 So the eqn is : x^2 + y^2 = 42200 ^2

OpenStudy (anonymous):

ty